Coastal Computer operates two retail outlets in Oakview, one on Main Street and the other in Lakeland Mall. The stores share the use of a central accounting department. The cost of the accounting department for last year was $166,000. Following are the operating results for the two stores for the year.
  

Main Street Lakeland Mall
Sales revenue $ 1,854,780 $ 3,026,220
Number of computers sold 2,331 1,369

  

Required:

a. Allocate the cost of the central accounting department to the two stores based on:

1. Number of computers sold.

2. Store revenue.
Answer #1

Answer:

1 Number of computers sold:
Main Street. 104,580 =166,000/(2331+1369)*2331
Lakwland mall 61,420. =166,000/(2331+1369)*1369
2. Store revenue:
Main street. 63,080 = 166,000/(18,54,780+30,26,220)*18,54,780
Lakeland mall 102,920 =166,000/(18,54,780+30,26,220)*30,26,220